### v2

# Tying knots in particle physics

Abstract
Lord Kelvin’s pioneering hypothesis that the identity of atoms is knots of vortices of the aether had a profound impact on the fields of mathematics and physics despite being subsequently refuted by experiments. While knot-like excitations emerge in various systems of condensed matter physics, the fundamental constituents of matter have been revealed to be elementary particles such as electrons and quarks, seemingly leaving no room for the appearance of knots in particle physics. Here, we show that knots indeed appear as meta-stable solitons in a realistic extension of the standard model of particle physics that provides the QCD axion and right-handed neutrinos. This result suggests that during the early Universe, a “knot dominated era” may have existed, where knots were a dominant component of the Universe, and this scenario can be tested through gravitational wave observations. Furthermore, we propose that the end of this era involves the collapse of the knots via quantum tunneling, leading to the generation of matter-antimatter asymmetry in the Universe. Our findings exhibit the significant role of knots in particle physics and represent a modern version of Kelvin’s hypothesis.
preprint: YGHP-24-06, DESY-24-104
Introduction
Knots, mathematically defined as closed curves embedded into three-dimensional space, appear not only when tying a tie but also in various scientific fields today, as pioneered by William Thomson, also known as Lord Kelvin. He conjectured that atoms, thought of as fundamental building blocks of matter at that time, are made from knots of vortices of the aether, which means that knots are the ultimate origins of all matter in our Universe Thomson:1869 . Although this elegant conjecture had attracted much attention from many physicists and mathematicians, it was eventually falsified by Michelson and Morley’s experiment ruling out the existence of aether, leading to the development of knot theory tait1898knots as a separate mathematical discipline.
Even though Kelvin’s idea failed to explain the most fundamental aspects of nature, the concept of knots has found extensive applications in various domains. One of the characteristics of knots is their topological nature; knots do not disappear unless the strings break or intersect, which has played a crucial role in persistent phenomena Kauffman:1991ds . For instance, knots appear as vortex knots or topological solitons called Hopfions Faddeev:1996zj in diverse subjects such as fluid dynamics Moffatt_1969 ; 10.1063/1.881574 ; Ricca:2009 ; Kleckner:2013 ; Arnold:2021 , superconductors Babaev:2001zy ; Rybakov:2018ktd , Bose-Einstein condensates PhysRevE.85.036306 ; Kleckner2016 ; Kawaguchi:2008xi ; Hall:2016 ; Ollikainen:2019dyh , 3He superfluids Volovik:1977 ; Volovik:2003fe , nematic liquid crystals PhysRevLett.110.237801 ; PhysRevLett.113.027801 ; Ackerman:2015 ; Ackerman:2017 ; Ackerman:2017b ; Tai:2018 ; Tai:2019 ; RevModPhys.84.497 ; Smalyukh:2020zin ; Smalyukh:2022 , colloids, magnets Kent:2020jvm , optics Dennis2010 , electromagnetism Trautman:1977im ; Ranada:1989wc ; Kedia:2013bw ; Arrayas:2017sfq , and active matter Shankar2022 . The stability of the solitons is ensured by the topological invariant associated with knots.
In modern physics, the understanding of the fundamental nature of matter has progressed from atoms to nucleons and ultimately to quarks and gluons primarily through elementary particle physics, instead of knots. Here a natural question may arise: Are knots truly irrelevant within the realm of elementary particle physics? It is known that a knot invariant in mathematics arises from topological field theories known as the Chern-Simons theory Witten:1988hf , which exhibits a direct connection between theoretical physics and mathematics. In addition, a knot as a soliton is proposed Faddeev:1996zj ; Battye:1998pe ; Manton:2004tk ; Radu:2008pp ; Kobayashi:2013xoa ; Shnir:2018yzp in a field-theoretic toy model called the Faddeev-Skyrme model, in which the stability of the knot soliton requires non-renormalizable interactions destroying the predictability of the model. Nevertheless, knots have received relatively less attention compared to condensed matter physics and other fields. In particular, knots have never appeared in phenomenologically viable models in the context of particle physics.
Here we revisit Kelvin’s idea in elementary particle physics. We theoretically show the existence of stable knot solitons in a particle physics model for the first time. The model is a simple setup containing global and local $U(1)$ symmetries denoted by $U(1)_{\mathrm{global}}$ and $U(1)_{\mathrm{local}}$, respectively. The two symmetries are spontaneously broken by condensations of two complex scalar fields, and thus a superfluid and superconductor are simultaneously realized, resulting in the production of two kinds of string-like defects called flux tubes and superfluid vortices, respectively. String loops (either knots or unknots) made of the single species of the flux tube or superfluid vortex shrink due to their tensions and eventually disappear PhysRevE.85.036306 ; Kleckner2016 . In contrast, it turns out that the knot solitons made by linking the two types of the strings remain stable. Such configurations are known as two-component links in knot theory. The stability is realized by the so-called Chern-Simons coupling between the Nambu-Goldstone (NG) boson and the massive $U(1)_{\mathrm{local}}$ gauge field. The gradient of the NG field and the magnetic flux of the gauge field induce a $U(1)_{\mathrm{local}}$ electric charge and hence electric flux around the string loop.111 Note that the stability mechanism by the Chern-Simon coupling is crucial, although knots were discussed in similar systems without such a coupling: liquid metalic hydrogen Babaev:2004rm ; Smiseth:2004na and the coexistence of a neutron superfluid and a proton superconductor inside neutron star interior Babaev:2002wa .
By identifying the $U(1)_{\mathrm{global}}$ and $U(1)_{\mathrm{local}}$ symmetries with the $U(1)$ Peccei-Quinn symmetry Peccei:1977hh ; Peccei:1977ur and $U(1)$ $B-L$ (baryon number minus lepton number) symmetry, which are popular hypothetical symmetries to resolve mysteries unanswered by the Standard Model (SM) of particle physics, our setup becomes a realistic model as a microscopic theory beyond the SM. In addition, we below argue a cosmological scenario that the knot solitons were produced at the early stage of the Universe. Because they can be sufficiently abundant to dominate the energy density of the Universe, this scenario is testable in terms of stochastic signals of gravitational waves. Furthermore, the knots eventually decay by a quantum effect, and the decay can produce an imbalance between matter and antimatter, which explains why matter is dominant compared to antimatter in the present Universe, see Fig. 1. Thus, this scenario is a modern revival of Kelvin’s dream; The Universe was full of knots, and matter originated from them.
Figure 1: Illustration of cosmic history in our scenario. After creation of the strings and knots, the knots eventually dominate the energy density of the Universe (knot dominated era). At $t\sim\tau_{\mathrm{decay}}$, they decay into light particles, leading to radiation dominated era again. This decay can produce the correct baryon asymmetry in the Universe via the non-thermal leptogenesis. In addition, the existence of the knot domination is imprinted on stochastic GW spectrum radiated from the string network, which can be tested by future GW observations.
The model
We start with a model that includes two complex scalar fields $\phi_{1}$ and $\phi_{2}$ and a $U(1)$ gauge field. The Lagrangian for the scalars is given as
$\mathcal{L}_{\phi}=|D_{\mu}\phi_{1}|^{2}+|D_{\mu}\phi_{2}|^{2}-V(\phi_{1},\phi_{2})$ (1)
with the potential term [^2] We here assume an exchange symmetry between the two scalars in the potential for simplicity. This is however not necessary in the following argument; we may add other terms such as $|\phi_{2}|^{4}$.
$V=\lambda(|\phi_{1}|^{2}+|\phi_{2}|^{2}-m^{2})^{2}-\kappa|\phi_{1}|^{2}|\phi_{2}|^{2}\,,$ (2)
and $\lambda,\kappa>0$. The Lagrangian is invariant under each $U(1)$ phase rotation of the scalars $\phi_{1}$ and $\phi_{2}$, which we denote $U(1)_{1}$ and $U(1)_{2}$, respectively. The covariant derivative is given as $D_{\mu}\phi_{1(2)}\equiv(\partial_{\mu}-iq_{1(2)}gA_{\mu})\phi_{1(2)}$, where $A_{\mu}$ is the $U(1)$ gauge field and $g$ is its coupling constant. $q_{1}$ and $q_{2}$ denote $U(1)$ gauge charges of $\phi_{1}$ and $\phi_{2}$, respectively. For simplicity, we take $q_{1}=1$ and $q_{2}=0$ for the moment, leading to that $U(1)_{1}$ is promoted to a gauge symmetry called $U(1)_{\mathrm{local}}$ while $U(1)_{2}$ is a global symmetry denoted by $U(1)_{\mathrm{global}}$. The Lagrangian for the kinetic term of $A_{\mu}$ is given as $\mathcal{L}_{g}=(-1/4)F_{\mu\nu}F^{\mu\nu}$ with $F_{\mu\nu}$ being the field strength. Thus the full Lagrangian is given as $\mathcal{L}=\mathcal{L}_{\phi}+\mathcal{L}_{g}$.
Because of the tachyonic mass parameter $-\lambda m^{2}<0$ in $V$, the two scalars condense and develop vacuum expectation values (VEVs) as $\langle\phi_{1}\rangle=\langle\phi_{2}\rangle=v/\sqrt{2}$, which spontaneously break both of the $U(1)_{\mathrm{local}}$ and $U(1)_{\mathrm{global}}$ symmetries.
At the broken phase, quantum corrections can induce the so-called Chern-Simons coupling in addition to the Lagrangian $\mathcal{L}_{\phi}+\mathcal{L}_{g}$, which characterizes the coupling between the NG boson $a\equiv-i\,\mathrm{arg}\,\phi_{2}$ and $A_{\mu}$, given as
$\mathcal{L}_{\mathrm{CS}}=C\,aF_{\mu\nu}\tilde{F}^{\mu\nu},$ (3)
where $\tilde{F}^{\mu\nu}=\epsilon^{\mu\nu\lambda\rho}F_{\lambda\rho}$ and $C$ is a model-dependent dimensionless constant; for instance, it may be induced by one-loop effects of fermions.333 If we assume the existence of a UV theory for our model, the coupling $C$ is quantized as $C=Ng^{2}/32\pi^{2}$ with an integer $N$. In such a case, this model possesses an emergent symmetry at low energy, characterized by a mathematical structure called the 4-group Hidaka:2021mml ; Hidaka:2021kkf while it is the 3-group in the unbroken phase of the $U(1)_{\mathrm{local}}$ symmetry Hidaka:2020izy ; Hidaka:2020iaz . To consider general setups, $C$ is taken as a free parameter in this paper. The model $\mathcal{L}_{\phi}+\mathcal{L}_{g}+\mathcal{L}_{\rm CS}$ includes the axion electrodynamics Wilczek:1987mv , which has been recently applied in condensed matter physics to describe topological superconductors Qi:2012cs ; Stone:2016pof ; Stalhammar:2021tcq .
The spontaneous breaking of the two symmetries $U(1)_{\mathrm{local}}$ and $U(1)_{\mathrm{global}}$ (or, $U(1)_{1}$ and $U(1)_{2}$) gives rise to two kinds of vortex strings as topological defects. We call these two strings the $\phi_{1}$ and $\phi_{2}$ strings; for the $\phi_{1}$ string, the phase of $\phi_{1}$ winds around the defect while for the $\phi_{2}$ string the phase of $\phi_{2}$ winds, both of whose windings are topologically protected. Since the $\phi_{1}$ string is a local string, it contains a magnetic flux associated with $A_{\mu}$, $|\Phi_{A}|=2\pi/g$, like flux tubes in superconductors while the $\phi_{2}$ string is a global string without magnetic fluxes like superfluid vortices.
Knot soliton
Figure 2: 3D plots of numerical solution for a knot soliton. The magenta and cyan regions represent the cores of the $\phi_{1}$ and $\phi_{2}$ strings ($|\phi_{1}|^{2}/v^{2}<0.1$ and $|\phi_{2}|^{2}/v^{2}<0.1$). The single $\phi_{2}$ string loop is linked with five $\phi_{1}$ string loops, so that the solution has the linking number $5$. In the middle and right panels, the arrows that are overlaid with transparent string cores indicate the magnetic flux $\vec{B}$ and the electric flux $\vec{E}$ of the massive gauge field $A_{\mu}$, respectively, and their colors correspond to their strength $|\vec{B}|$ and $|\vec{E}|$. The magnetic flux is along the $\phi_{1}$ strings, one of whose direction is indicated by the large gray arrow, while the electric one comes out from the strings, which means that the strings have net electric charges. We here take $\lambda/g^{2}=10^{3}$, $\kappa/g^{2}=0.0008$, and $C=400$. The energy of the solution is $7.0\times 10^{3}\,v/g$.
Figure 3: 3D plots of numerical solutions for other knot solitons. The colors and the parameters are the same as those in Fig. 2. The two left plots have the same linking number $4$ and the the other plots have the linking number $5$. The energy of the solutions are $6.0\times 10^{3}\,v/g$, $6.3\times 10^{3}\,v/g$, $7.3\times 10^{3}\,v/g$ and $7.5\times 10^{3}\,v/g$ (from left to right).
The Chern-Simons coupling (3) does not affect the individual $\phi_{1}$ or $\phi_{2}$ string because the $\phi_{1}$ ($\phi_{2}$) string does not contain excitation of the NG boson $a$ (the gauge field $A_{\mu}$), resulting in that (3) vanishes. Thus a loop made of either the $\phi_{1}$ or $\phi_{2}$ string is not stable but shrinks due to their tension and eventually disappears.
Now, make the $\phi_{1}$ and $\phi_{2}$ strings linked. Particularly, one simple example is a link made of a single $\phi_{2}$ string and multiple $\phi_{1}$ strings (See the left panel of Fig. 2). Now the coupling (3) plays a crucial role; it induces the electric flux of the $A_{\mu}$ field around the $\phi_{1}$ string. To see this, let us take a functional derivative of the Chern-Simons coupling with respect to $A_{0}$ as
$\frac{\delta}{\delta A_{0}}\int d^{4}x\,\mathcal{L}_{\mathrm{CS}}$ $=-C\,\left(\vec{\nabla}a\cdot\vec{B}\right)\,,$ (4)
where $\vec{B}$ is the magnetic field of $A_{\mu}$, $\vec{B}=\nabla\times\vec{A}$, and we have used that all fields are static. Eq. (4) changes the Gauss law in the Maxwell equation with respect to $A_{0}$ and hence $C\,\vec{\nabla}a\cdot\vec{B}$ behaves as a source of the electric field. 444 A similar effect is well known for Chern-Simons vortices in $2+1$ dimensions Horvathy:2007ps ; Horvathy:2008hd . The total electric charge is estimated to be $C\int d^{3}x\,\vec{\nabla}a\cdot\vec{B}=4\pi^{2}\,CN_{\mathrm{link}}/g$ where the strings are assumed to be sufficiently thin Yokoi_2013 . The integer $N_{\mathrm{link}}$ is the linking number of the $\phi_{1}$ and the $\phi_{2}$ strings. [^5] Formally this is the same as cross helicity in magnetohydro dynamics when we regard $\vec{\nabla}a$ as the velocity of fluid Yokoi_2013 ; Nastase:2022aps . From this, the $\phi_{1}$ strings linking with the $\phi_{2}$ strings get the electric charge, which gives rise to the electric field $\vec{E}$ around the strings. Even though this electric field is screened due to the VEV of $\phi_{1}$ like Meissner effect, for large enough $C$, the $\phi_{1}$ string loops are prevented from shrinking by the electric field and is stabilized.666 A similar vortex loop with the electric charge/current is known as the vorton Davis:1988ij ; Davis:1988jq . The classical and quantum studies of the stability of vortons are given in Refs. Battye:2008mm ; Radu:2008pp ; Garaud:2013iba ; Battye:2021sji ; Battye:2021kbd ; Ibe:2021ctf ; Abe:2022rrh . Our knot soliton considered here is however different from those in the sense that it does not contain internal degrees of freedom on the loop.
In order for the linked configuration to be stable under arbitrary perturbations, another condition $\lambda\gg g^{2},\kappa$ is necessary. If this is met, the link of the $\phi_{1}$ and $\phi_{2}$ strings cannot be de-linked, i.e., the strings cannot pass through each other. This is because in the limit of $\lambda\to\infty$, the linking number is equivalent to the Skyrmion number Gudnason:2020luj ; Gudnason:2020qkd which is characterized by the homotopy group $\pi_{3}(S^{3})\simeq{\mathbb{Z}}$ defined by a map from $S^{3}$, obtained by a compactification of the real space ${\mathbb{R}}^{3}$, to $S^{3}$ defined by $|\phi_{1}|^{2}+|\phi_{2}|^{2}=\mathrm{const.}$, and hence in this regime, this number approximately becomes a topological invariant.777 This is similar to the case of Skyrmions in two-component Bose-Einstein condensates Ruostekoski:2001fc ; Battye:2001ec ; Nitta:2012hy . In practical cases with finite $\lambda$, the $\phi_{1}$ and $\phi_{2}$ strings are repelled from each other Eto:2011wp ; Rybakov:2018ktd as far as the condition $\lambda\gg g^{2},\kappa$ is met, and hence they cannot be overlapped.
We numerically solve the classical equations of motion (EOMs) to confirm the existence of the stable knot soliton in the model. We perform a minimization procedure for the energy of the configuration starting from an appropriate initial configuration with the Gauss-law constraint. See Appendix .1 for the details. After the minimization converges with desired accuracy, the configuration is regarded as the static solution of the EOMs. Fig. 2 shows the cores of the strings of the obtained solution, in which the magenta and cyan colors represent regions $|\phi_{1}/v|^{2}<0.1$ and $|\phi_{2}/v|^{2}<0.1$, respectively. We here take $\lambda/g^{2}=10^{3}$, $\kappa/g^{2}=0.0008$, and $C=400$. The magnetic and electric fields are also shown by arrows. In this figure, the single $\phi_{2}$ string loop is linked with five $\phi_{1}$ string loops (with the same linking direction), so that the solution has the linking number $5$. One can also see that the magnetic flux is along the $\phi_{1}$ strings, one of whose direction is indicated by the large gray arrow. On the other hand, the electric one comes out from the strings, which means that the strings have net electric charges. This is consistent with the argument above. 888 It is known that the electric flux may cause an instability for the NG boson and gauge field due to the Chern-Simons coupling Ooguri:2011aa . Even though our setup has some differences from theirs, such as the mass of the gauge field, this effect still exists in our case. Consequently, the NG boson $a$ and gauge field $A_{i}$ exhibit oscillating profiles inside the $\phi_{1}$ strings. However, profiles inside the strings are not crucial for the stability because the amount of the electric charge is fixed by the linking number unless the strings do not have overlap.
Note that a similar solution is obtained by acting parity transformation on the knot soliton shown here. This solution has the opposite linking direction of the $\phi_{1}$ and $\phi_{2}$ strings compared to the knot soliton, so that we call this the anti-knot soliton. The anti-knot one has the $U(1)_{X}$ electric charge (and Skyrme number) with the opposite signs to those of the knot one.
We can also obtain other knot solitons, which are shown in Fig. 3. The leftmost panel shows a single loop of the $\phi_{2}$ string linking with four $\phi_{1}$ string loops. In the other panels, single $\phi_{1}$ and $\phi_{2}$ loops make higher linking number by linking multiple times. The colors are the same as those in Fig. 2. For both types, the energy grows as the linking number gets higher because the electric charge contained in the $\phi_{1}$ string becomes larger. The energy of all configurations are almost dominated by the $\phi_{1}$ strings and the electric flux. Within the same linking number, the energy gets larger as the length of the $\phi_{1}$ string gets longer. Therefore, the leftmost solution in Fig. 3 and that in Fig. 2 have the least energy with the linking number $4$ and $5$, respectively. Although we did not find any more other knot solitons with our numerical code, this does not necessarily mean non-existence of other solitons.
Even when the (anti-)knot solitons are classically stable, they can decay due to quantum tunneling, whose lifetime is denoted by $\tau_{\mathrm{decay}}$. One possible decay process is the de-linking process, i.e., the $\phi_{1}$ and $\phi_{2}$ strings pass through each other by quantum tunneling. The calculation of the decay rate is complicated and beyond the scope of this paper. Instead, we will later relate the decay rate to the cosmological history of the knot solitons.
Particle phenomenology
In spite of the great success of the SM, there are still several problems that the SM is not able to answer, such as dark matter of the Universe, tiny masses of the neutrinos, strong CP problem, and baryon asymmetry in the Universe. (For a review, see, e.g., Ref. ParticleDataGroup:2022pth .) It is believed that (some of) these problems should be resolved by new physics beyond the SM.
Particularly, one elegant way to extend the SM is to introduce new symmetries that do not exist within the SM. The strong CP problem and dark matter are simultaneously explained by introducing a global $U(1)$ symmetry called the Peccei-Quinn symmetry Peccei:1977hh ; Peccei:1977ur (denoted by $U(1)_{\mathrm{PQ}}$), which spontaneously breaks and provides a NG boson, the QCD axion Weinberg:1977ma ; Wilczek:1977pj . In addition, it is well motivated to introduce the gauge $B-L$ (baryon number minus lepton number) symmetry, $U(1)_{B-L}$, [^9] The SM contains global $U(1)_{B-L}$ symmetry as an accidental symmetry although it cannot be gauged due to the gauge anomaly in the SM unless right-handed neutrinos are introduced. because it requires the existence of the right-handed neutrinos and its spontaneous breaking gives them Majorana masses, which naturally explain the tiny neutrino masses in terms of the type-I seesaw mechanism  Minkowski:1977sc ; Gell-Mann:1979vob ; Mohapatra:1979ia ; Yanagida:1979as . The right-handed neutrinos are also expected to play key roles to generate baryon asymmetry in the Universe. Furthermore, the $U(1)_{B-L}$ gauge symmetry naturally arises from grand unified theories (GUTs).
Now we apply our study on the knot soliton to particle physics models and discuss its influence on cosmology. One realistic and promising setup is obtained by identifying the $U(1)_{\mathrm{local}}$ and $U(1)_{\mathrm{global}}$ symmetries in (1) as the $U(1)_{B-L}$ and $U(1)_{\mathrm{PQ}}$ symmetries, respectively. In order for $\phi_{1}$ to couple with right-handed neutrinos, the $B-L$ charge of $\phi_{1}$ is taken as $q_{1}=2$ instead of $1$ while keeping $q_{2}=0$. We emphasize that this setup is well motivated because it can elegantly resolve some of the mysteries in particle physics as stated above.101010 When one allows $\phi_{2}$ to have a small non-zero charge under $U(1)_{B-L}$ and introduces fermionic matters charged under the $U(1)_{B-L}$ and $U(1)_{\mathrm{PQ}}$ symmetries, the setup may have a bonus Fukuda:2017ylt ; Ibe:2018hir that the quality of the $U(1)_{\mathrm{PQ}}$ symmetry is protected against the effect of quantum gravity. (See Refs. Fukuda:2017ylt ; Ibe:2018hir for more details.) In this setup, the QCD axion is identified with the phase of $\phi_{2}$. For this to solve the strong CP problem, we need additional fermionic fields inducing the Chern-Simons coupling between the QCD axion and gluon, which may be realized by introducing Kim-Shifman-Vainshtein-Zakharov(KSVZ)-like Kim:1979if ; Shifman:1979if heavy quarks. How the QCD axion couples to the $B-L$ and SM gauge bosons depends on the charge assignment of the fermionic sectors, which is kept as general as possible hereafter. From the astrophysical constraints on the QCD axion, we take $v\gtrsim 10^{8}\,\mathrm{GeV}$ Irastorza:2018dyq ; Chang:2018rso ; Hamaguchi:2018oqw ; Caputo:2024oqc .
Cosmology and gravitational wave
Now let us discuss cosmological fate of the knot solitons. The Universe was very hot and full of thermal relativistic particles at early stages (after the cosmic inflation). When the temperature $T$ of the Universe is much larger than $v$, the $U(1)_{B-L}$ and $U(1)_{\mathrm{PQ}}$ symmetries are not broken while they get spontaneously broken almost simultaneously due to a phase transition when $T$ gets less than a critical temperature $T_{c}\simeq v$. At the transition, the $\phi_{1}$ and $\phi_{2}$ strings are produced by the Kibble-Zurek mechanism Kibble:1980mv ; Zurek:1985qw . The produced strings have random configurations, in which some of the strings are linked with finite probability. After the oscillation and the random motion are relaxed, long $\phi_{1}$ and $\phi_{2}$ strings form conventional networks whose typical length scale is the Hubble scale. Besides them, small loops with the linking configurations become the knot solitons that we have studied above. Let us estimate the number density of the knot solitons produced by this mechanism. One should note that, when $\kappa$ is sufficiently small compared to the other parameters $\lambda,g$ and $C$, the scalar potential $V$ has an approximate global $O(4)$ symmetry spontaneously broken down to $O(3)$. Thus, within a microscopic length scale smaller than $1/(\sqrt{\kappa}T_{c})$, the linking number density is well approximated by the number density of textures Turok:1989ai ; Borrill:1991mv ; Leese:1991gt ; Vilenkin:2000jqa rather than that of linking cosmic strings Vachaspati:1994ng . According to Ref. Turok:1989ai , the linking number density is given as $\sim 0.04\,\xi^{-3}$, where $\xi$ is the correlation length to be $\xi\sim T_{c}^{-1}$. Noting that the knot solitons that we found above should have the linking number larger than or equal to $4$, the number density of the (anti-)knot solitons is estimated as $n_{\mathrm{knot}}(T=T_{c})\sim(0.04)^{4}\,\xi^{-3}$. [^11] Note that results given below do not depend on the value of the produced number density $n_{\mathrm{knot}}(T_{c})$ as long as they dominate the energy density of the Universe before they decay.
After the production, the knot solitons behave as heavy particles decoupled from the thermal bath with the long lifetime $\tau_{\mathrm{decay}}$, and hence their energy density is given as $\rho_{\mathrm{knot}}=M\,n_{\mathrm{knot}}$ with the mass of the knots $M\simeq\mathcal{O}(10^{3})\,v/g$. Since $\rho_{\mathrm{knot}}$ is diluted by the cosmic expansion being proportional to the cubic of the scale factor $a(t)^{-3}$, which is slow compared to the radiation energy density in the thermal plasma, $\rho_{\mathrm{rad}}\propto a(t)^{-4}$, the energy density of the knots eventually overcomes the latter and dominates the total energy density of the Universe when the temperature reaches $T\sim M\,n_{\mathrm{knot}}(T_{c})/(g_{*}T_{c}^{3})$. Here $g_{*}$ is the number of the degrees of freedom for relativistic particles, $g_{*}=\mathcal{O}(10^{2})$. Note that this domination persists until they decay. We call this epoch “knot dominated era”. When cosmic time exceeds the lifetime of the knots, $t\sim\tau_{\mathrm{decay}}$, they decay by the quantum tunneling releasing their energy into light particles, which leads to the radiation dominated epoch again via a secondary reheating. Assuming instantaneous decays, the temperature of the reheated thermal bath is found to be Kolb:1990vq
$T_{rh}$ $\simeq\left(\frac{30}{\pi^{2}g_{*}}\frac{3M_{\mathrm{pl}}^{2}}{8\pi}\,\Gamma^{2}\right)^{1/4}\,.$ (5)
with $\Gamma\equiv 1/\tau_{\mathrm{decay}}$ and $M_{\mathrm{pl}}$ being the decay rate and the Planck mass, respectively. Note that $\Gamma$ is typically exponentially suppressed because of non-perturbative processes. In order to realize the successful Big-Bang Nucleosynthesis (BBN), $T_{rh}$ must be larger than $\mathcal{O}(1)$ MeV Kawasaki:1999na ; Kawasaki:2000en ; Hannestad:2004px . Therefore, the Universe experiences the knot dominated era sandwiched by two radiation dominated eras before BBN. See Fig. 1.
Remarkably, this existence of the knot dominated era can be observed in terms of gravitational waves (GW) when there exists a sufficient GW source. Fortunately, this model already contains a possible source of GW: the string network. Thermal history of the Universe including the knot dominated era is imprinted on the GW spectrum. Following the study given in Refs. Cui:2017ufi ; Cui:2018rwi ; Gouttenoire:2019kij , we obtain an expected spectrum of the stochastic GW emitted from the network of the $\phi_{1}$ strings depending on $v$ and $T_{rh}$ as shown in Fig. 4. Here $\Omega_{\mathrm{GW}}h^{2}$ is defined as $\Omega_{\mathrm{GW}}h^{2}=(d\rho_{\mathrm{GW}}(f)/d\log f)h^{2}/\rho_{c}$ with $h$ being the dimensionless Hubble constant, $\rho_{\mathrm{GW}}(f)$ the GW energy density for a frequency $f$ observed today, and $\rho_{c}$ the critical energy density. We also show in Fig. 4 the observed signals of NANOGrav 15-year data set NANOGrav:2023gor as gray region 121212 We do not aim to explain the NANOGrav 15-yrs signals by the GW spectrum from the $\phi_{1}$ strings because the signals are unlikely to be explained by cosmic strings alone NANOGrav:2023hvm . Instead they may be explained by combining the spectrum of the strings with those of super massive black holes. and projected sensitivity curves of future GW observatories. The data points of the projected sensitivity curves are taken from Ref. Schmitz:2020syl . One can see that the spectra are deviated due to the existence of the knot dominated era (solid lines) and can be distinguished from the conventional spectra without the knot domination (red dashed lines).
Figure 4: Stochastic GW spectrum $\Omega_{\mathrm{GW}}h^{2}$ from the network of the $\phi_{1}$ strings with $G\mu=10^{-11}$, in which $G$ and $\mu$ are the Newton constant and the $\phi_{1}$ string tension given as $\mu\simeq v^{2}$. The red dashed line corresponds to the cosmological evolution without the knot solitons while the solid ones include the knot dominated era with different reheating temperatures due to the knot decay, $T_{rh}=5$ MeV (blue), $100$ GeV (orange), and $10^{5}$ GeV (green). We also show the observed signals of NANOGrav 15-year data set NANOGrav:2023gor as gray region and projected sensitivity curves of future observatories: DECIGO Kawamura:2020pcg (orange), Cosmic Explorer (CE) Reitze:2019iox (green), LISA Bartolo:2016ami (blue), and Square Kilometer Array (SKA) Janssen:2014dka (red).
Baryogenesis
Our Universe is baryon asymmetric, i.e., baryon is more than anti-baryon as characterized by a quantity $Y_{B}\equiv(n_{B}-n_{\bar{B}})/s\simeq 0.8\times 10^{-10}$ Planck:2018vyg where $n_{B(\bar{B})}$ and $s$ are the number density of (anti)baryon and the entropy density, respectively. The SM however cannot explain this asymmetry. As stated above, the knot solitons remain in the Universe and later decay into particles. When $T_{rh}$ is larger than the electroweak (EW) scale $\mathcal{O}(10^{2})$ GeV, 131313 There can be an alternative scenario of baryogenesis even when $T_{rh}$ is smaller than the electroweak scale. This is complementary to the scenario presented here. See Appendix .3 for the details. this decay process can produce baryon asymmetry in the Universe in a similar way to non-thermal leptogenesis via inflaton decay Asaka:1999yd ; Giudice:1999fb ; Asaka:1999jb ; Lazarides:1990huy ; Kumekawa:1994gx . Note that the baryon asymmetry produced at earlier stages than the end of the knot domination is diluted by the secondary reheating due to the knot decay. For simplicity, we here assume it to be sufficiently diluted away and concentrate on the baryon asymmetry produced by the knot decay.
To see that this scenario works, we should note that to ensure cancellation of the gauge anomaly of $U(1)_{B-L}$, this model contains the right-handed (RH) neutrinos $\nu_{R}$ described by the Lagrangian
[…]
where $H$ and $\ell$ are the SM Higgs doublet and lepton doublet, respectively. $y_{R}^{ij}$ ($y_{D}^{ij}$) is the Majorana-type (Dirac-type) Yukawa coupling constant. The indices $i,j$ denote the generations and run as $i,j=1,2,3$. After $\phi_{1}$ takes the VEV $v$, the RH neutrinos get Majorana masses, whose mass eigenvalues and eigenstates are denoted by $M_{Ri}$ and $N_{i}$, respectively. We suppose $M_{R1}<M_{R2}<M_{R3}$, so that the baryon asymmetry is dominantly produced via the lightest state $N_{1}$. Through the couplings with $\phi_{1}$ and $A_{\mu}$ in $\mathcal{L}_{R}$, the decay of a knot soliton produces $N_{i}$. We parameterize an energy fraction of the produced $N_{1}$ compared to the knot static energy as $f_{N_{1}}\in[0,1]$, i.e., the produced number of $N_{1}$ is given as $f_{N_{1}}M/M_{R1}$, in which $f_{N_{1}}$ is not expected to be very small when $M_{R1}$ is smaller than the gauge boson mass.
Typically $N_{1}$ has much larger decay rate, whose main decay channels are $N_{1}\to H+\ell$ and $N_{1}\to H^{\dagger}+\bar{\ell}$, than that of the knot soliton, so that the produced $N_{1}$ immediately decay into $H$ ($H^{\dagger}$) and $\ell$ ($\bar{\ell}$). When their decay rates are different, the lepton asymmetry $n_{L}-n_{\bar{L}}$ can be produced. See Appendix .2 for the calculation of the lepton asymmetry. Through the EW sphaleron process as usual, some of the lepton asymmetry is converted into the baryon asymmetry. Thus leptogenesis is followed by baryogenesis.
When the masses $M_{R1},M_{R2},M_{R3}$ are not degenerated, the produced baryon asymmetry is maximized for a hierarchical case $M_{R1}\ll M_{R2},M_{R3}$ Flanz:1994yx ; Covi:1996wh ; Buchmuller:1997yu ; Davidson:2002qv , leading to
$Y_{B}$ $\simeq 4.1\times 10^{-11}\,f_{N_{1}}\left(\frac{T_{rh}}{10^{6}\,\mathrm{GeV}}\right)\left(\frac{m_{3}}{0.05\,\mathrm{eV}}\right)\delta_{\mathrm{eff}}\,,$ (7)
where $m_{3}$ and $\delta_{\mathrm{eff}}$ are the heaviest mass of the active neutrinos and the effective CP phase coming from the Yukawa matrices, respectively. If one takes $f_{N}=\delta_{\mathrm{eff}}=\mathcal{O}(1)$, the correct baryon asymmetry can be explained for $T_{rh}\simeq 10^{6}$ GeV. This case is however difficult to test in terms of GW spectrum by near future experiments because it requires sensitivity for higher frequency region than $10^{5}$ Hz as shown in Fig. 4.
On the other hand, when the masses of $N_{1}$ and $N_{2}$ are almost degenerated, $Y_{B}$ can be enhanced to be as large as
$Y_{B}$ $\simeq 0.8\times f_{N_{1}}\frac{T_{rh}}{M_{R1}}$ (8)
like resonant leptogenesis Pilaftsis:1997jf ; Pilaftsis:2003gt . In such a case, the reheating temperature $T_{rh}$ can be lowered to, say, $\mathcal{O}(10^{2})$ GeV for $M_{R1}=10^{12}$ GeV. This case is quite attractive because the stochastic GW spectrum lies within the range of CE, and hence this scenario is expected to be tested. (See the orange line in Fig. 4.)
To summarize, within reasonable setups, the decay of the knot solitons can produce the correct baryon asymmetry, which gives the matter-antimatter asymmetry in the present Universe. In this sense, although the knot itself is not the building block of matter as conjectured by Kelvin, it might be the origin of matter and a key ingredient of nature. Fig. 1 shows the summary of cosmological history.
For further studies in the phenomenological perspective, we need concrete model building. Particularly, it is worth considering embedding this setup into GUTs such as $SO(10)$ GUTs since $U(1)_{B-L}$ and the SM gauge group are elegantly embedded into $SO(10)$ as is well-known. Although we have taken $C$, the coefficient of the Chern-Simons coupling, as a free parameter, it may be explained as one-loop effect of matter fermions charged under the $U(1)_{B-L}$ and $U(1)_{\mathrm{PQ}}$ symmetries. It is not obvious whether the parameter space and field contents that ensure the existence of the knot soliton is compatible with GUTs. If so, the knot soliton can be regarded as an indirect probe of GUTs. In any case, the knot soliton in particle physics has many potential applications to understand fundamental aspects of nature.
Acknowledgments
The authors would like to thank Simone Blasi, Asuka Ito, Ryusuke Jinno, Kohei Kamada, Yuka Kotorii, Shunya Mizoguchi, Nobuyuki Matsumoto, Hiroto Shibuya, Fumio Uchida, Wen Yin and Ryo Yokokura for useful comments. The numerical calculations were carried out on Yukawa-21 at YITP in Kyoto University. This work is supported in part by JSPS Grant-in-Aid for Scientific Research KAKENHI Grant No. JP22H01221 (M. E. and M. N.) and No. JP22KJ3123 (Y. H.) and by the Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy - EXC 2121 Quantum Universe - 390833306. The work of M. N. is supported in part by the WPI program “Sustainability with Knotted Chiral Meta Matter (SKCM2)” at Hiroshima University. The photo of a galaxy used in Fig. 1 is taken from NASA Image and Video Library.
Appendix
.1 Numerical technique to obtain knot solitons
We give numerical technique performed in this work. Because we are interested in static solutions of the EOMs, it is sufficient to minimize the energy functional,
$E$ $=\int d^{3}x\,\mathcal{E}$ (9)
with
$\mathcal{E}=$ $(D_{i}\phi_{1})^{2}+(D_{i}\phi_{2})^{2}+V(\phi_{1},\phi_{2})+\frac{1}{2}(\partial_{i}A_{j})^{2}$
$-\frac{1}{2}(\partial_{i}A_{0})^{2}-gA_{0}J_{0}+C\epsilon^{ijk}\partial_{i}aA_{0}\partial_{j}A_{k}$
$+\frac{\gamma-1}{2}(\partial_{i}A_{i})^{2}\,,$ (10)
and
$J_{0}\equiv gA_{0}(q_{1}^{2}|\phi_{1}|^{2}+q_{2}^{2}|\phi_{2}|^{2})\,,$ (11)
where we have added the gauge fixing term $\gamma(\partial_{i}A_{i})^{2}/2$ to realize the Coulomb gauge $\partial_{i}A_{i}=0$ and $\gamma$ is the gauge fixing parameter.
We here meet two obstructions to perform the minimization.
  • The energy density (10) is apparently not positive definite because the temporal component $A_{0}$ has the “wrong sign”.
  • For large $C$, the first-order differential term $A_{0}\partial_{j}A_{k}$ is dominant over the kinetic term $(\partial_{i}A_{j})^{2}$.
Due to the first obstruction, a naive minimization with respect to $A_{0}$ does not converge. To avoid this, we solve separately the EOM of $A_{0}$ called the Gauss-law constraint,
$0=\frac{\delta E}{\delta A_{0}}=\partial^{2}A_{0}-2gJ_{0}+C\epsilon^{ijk}\partial_{i}a\partial_{j}A_{k}\,,$ (12)
from which it follows that Eq. (10) is rewritten as
$\mathcal{E}=$ $(D_{i}\phi_{1})^{2}+(D_{i}\phi_{2})^{2}+V(\phi_{1},\phi_{2})+\frac{1}{2}(\partial_{i}A_{j})^{2}$
$+\frac{1}{2}C\rho A_{0}+\frac{\gamma-1}{2}(\partial_{i}A_{i})^{2}$ (13)
with $\rho=\epsilon^{ijk}(\partial_{i}a)\partial_{j}A_{k}$. [^14] This expression can be made obviously positive definite when one uses the following formal solution of the Gauss law, $A_{0}=C\left[-\partial_{i}^{2}+2g^{2}(q_{1}^{2}|\phi_{2}|^{2}+q_{2}^{2}|\phi_{2}|^{2})\right]^{-1}\epsilon^{ijk}\partial_{i}a\partial_{j}A_{k},$ from which the fifth term in Eq. (13) becomes $\frac{C^{2}}{2}\rho\,\left[-\partial_{i}^{2}+2g^{2}(q_{1}^{2}|\phi_{2}|^{2}+q_{2}^{2}|\phi_{2}|^{2})\right]^{-1}\,\rho\,.$
The second point often produces a numerical instability when we discretize the system. To resolve this, we introduce an auxiliary field $B_{i}$ and replace $\rho$ as
$\rho\to B_{i}\partial_{i}a$ (14)
in Eq. (13) imposing constraints
$B^{i}=\epsilon^{ijk}\partial_{j}A_{k}\quad(i=1,2,3)\,.$ (15)
Thanks to this, there is no first-order differential term apparently. In order to realize these constraints, we follow the procedure presented in Ref. Hestenes1969MultiplierAG . We introduce multiplier fields $w_{i}(x)$ and add terms
$w_{i}(B^{i}-\epsilon^{ijk}\partial_{j}A_{k})+\frac{U}{2}(B^{i}-\epsilon^{ijk}\partial_{j}A_{k})^{2}$ (16)
to the energy density ($U=\mathrm{const.}$). It is shown in Ref. Hestenes1969MultiplierAG that these constraints are achieved by updating $w_{i}$ iteratively as follows
$w_{i}\to w_{i}+U(B^{i}-\epsilon^{ijk}\partial_{j}A_{k})$ (17)
when $U$ is sufficiently large.
Then the task is to minimize the three-dimensional integration of Eq. (13) in which $A_{0}$ is determined by solving the Gauss law constraint (12) and the constraints (15) are imposed. We introduce a naive spacial discretization with lattice spacing $d$ and approximate the derivatives by second-order central-difference scheme. Denoting the all variables on the lattice sites by $u_{a,j}=\{\mathrm{Re}\,\phi_{1,2}(x_{j}),\mathrm{Im}\,\phi_{1,2}(x_{j}),A_{i}(x_{j}),B_{i}(x_{j})\}$ and $s_{j}=A_{0}(x_{j})$, we have $E\simeq d^{3}\sum\mathcal{E}_{\mathrm{disc}}(u_{a,j},s_{j})$, where the summation is taken over all lattice points.
We adopt the so-called relaxation procedure, in which one starts from appropriate values $u_{a}^{(0)}$ and solve an iterative equation,
$u_{a,j}^{(n+1)}=u_{a,j}^{(n)}-\alpha\frac{\partial\mathcal{E}_{\mathrm{disc}}(u^{(n)},s^{(n)})}{\partial u_{a,j}^{(n)}}\,,$ (18)
to obtain $(n+1)$-th step values from $n$-th step ones. Here, $\alpha$ is the step size. To determine $s_{j}^{(n+1)}$, we solve the Gauss law numerically at each step by the similar procedure,
$s_{j}^{(n+1)}$ $=s_{j}^{(n)}$
$+\beta\left[\Delta s_{j}^{(n)}-2gJ_{0}(u_{a,j}^{(n+1)},s_{j}^{(n)})+C\rho(u_{a,j}^{(n+1)},s_{j}^{(n)})\right]$ (19)
where $\Delta$ is the discretized Laplacian and $\beta$ is the step size for $s_{j}$. In addition, we have to repeat the update for the multiplier $w_{i}$ given by Eq. (17). To summarize, we repeat the following procedure:
[…]
If this iteration converges, the configuration is the static solution of the EOM with the conditions (12) and (15).
In this work, we take $\gamma=1+U$, $U=50$, $d=0.8/v$, $\alpha=4\times 10^{-4}v^{-2}$, and $\beta=2\times 10^{-3}v^{-2}$. The number of lattice sites is $320^{3}$. We use a convergence criteria that
$\sum_{a}\overline{\left(u_{a,j}^{(n+1)}-u_{a,j}^{(n)}\right)^{2}}<1.06\times 320^{-3}\,,$ (20)
where the overline indicates the spatial average.
.2 Calculation of non-thermal leptogenesis
We here give more detailed calculation for leptogenesis from the knot decay. After $N_{1}$’s are produced from the knots, they further decay into $H$ ($H^{\dagger}$) and $\ell$ ($\bar{\ell}$). It is convenient to introduce the so-called lepton asymmetry parameter defined through decay rates of the processes as
$\epsilon\equiv\frac{\Gamma(N_{1}\to H+\ell)-\Gamma(N_{1}\to H^{\dagger}+\bar{\ell})}{\Gamma(N_{1}\to H+\ell)+\Gamma(N_{1}\to H^{\dagger}+\bar{\ell})}\,,$ (21)
which depends on parameters of the model. Then, the produced lepton asymmetry per knot soliton is expressed as $\epsilon f_{N_{1}}M/M_{R1}$.
Since $T_{rh}$ is assumed to be larger than the EW scale, the EW sphaleron process is equilibrated and convert the lepton number into the baryon number as
$Y_{B}$ $=-\left.\frac{28}{79}\frac{n_{L}-n_{\bar{L}}}{s}\right|_{\mathrm{decay}}$
$\simeq-\left.\frac{28}{79}\frac{\epsilon f_{N_{1}}n_{\mathrm{knot}}M}{M_{R1}}\frac{45}{2\pi}\frac{1}{g_{s}T_{rh}^{3}}\right|_{\mathrm{decay}}$
$\simeq-\frac{21\pi}{79}\epsilon f_{N_{1}}\frac{T_{rh}}{M_{R1}}\,,$ (22)
where $g_{s}(\simeq g_{*})$ is the effective relativistic degrees of freedom for entropy. Note that $M_{R1}$ should satisfy $M_{R1}>T_{rh}$ to avoid the wash-out effect.
When the masses $M_{R1},M_{R2},M_{R3}$ are not degenerated, the parameter $\epsilon$ is maximized for a hierarchical case $M_{R1}\ll M_{R2},M_{R3}$, as given by Flanz:1994yx ; Covi:1996wh ; Buchmuller:1997yu ; Davidson:2002qv
$\epsilon\simeq-\frac{3}{8\pi}\frac{M_{R1}}{\langle H\rangle^{2}}m_{3}\delta_{\mathrm{eff}}\,.$ (23)
Substituting Eq. (23) into Eq. (22), one gets Eq. (7).
On the other hand, when the masses of $N_{1}$ and $N_{2}$ are almost degenerated, $\epsilon$ can be enhanced to be as large as $\mathcal{O}(1)$ like resonant leptogenesis Pilaftsis:1997jf ; Pilaftsis:2003gt . In such a case, $Y_{B}$ is given by Eq. (8).
.3 Baryogenesis via linking flux
We present an alternative scenario that the knot decay may produce baryon asymmetry in the Universe even when $T_{rh}$ is smaller than the EW scale $\mathcal{O}(10^{2})$ GeV. The key is the fact that the $\phi_{2}$ string also has the magnetic flux of the $U(1)_{B-L}$ gauge field and becomes a fractional vortex when $\phi_{2}$ has a non-zero $B-L$ charge, $q_{2}\neq 0$. This allows the knot solitons to have the linking magnetic fluxes of the $U(1)_{B-L}$ gauge field. This linking flux has the Chern-Simons number (or helicity) $N_{CS}[A]\sim q_{1}q_{2}N_{\mathrm{link}}$.
In general, this $U(1)_{B-L}$ gauge field $A_{\mu}$ has a kinetic mixing with the $U(1)_{Y}$ gauge field in the SM, $\mathcal{L}_{\mathrm{mix}}=(\epsilon_{\mathrm{mix}}/2)F_{\mu\nu}Y^{\mu\nu}$, especially when they arise from a unified simple gauge group such as $SO(10)$. Due to the kinetic mixing, the magnetic flux of $A_{\mu}$ induces the flux of $Y_{\mu}$ along the $\phi_{1}$ and $\phi_{2}$ strings. Therefore a knot soliton with the Chern-Simons number of $U(1)_{X}$ must contain the Chern-Simons number of $U(1)_{Y}$, $N_{CS}[Y]\sim\epsilon_{\mathrm{mix}}^{2}N_{CS}[A]\sim\epsilon_{\mathrm{mix}}^{2}q_{1}q_{2}N_{\mathrm{link}}$. In the broken phase of the EW symmetry, the $U(1)_{Y}$ flux is dressed as the $Z$ boson flux Hyde:2013fia , which is a superposition of the $U(1)_{Y}$ and $SU(2)_{W}$ fluxes. (The unbroken $U(1)$ electromagnetic flux is not induced.)
When the knots decay at cosmic time $t\simeq\tau_{\mathrm{decay}}$, all the magnetic fluxes contained in the knots disappear, and hence $N_{CS}[Y]$ changes from $\epsilon_{\mathrm{mix}}^{2}q_{1}q_{2}N_{\mathrm{link}}$ into 00 per knot. This decay process produces the baryon number because the baryon number is not conserved due to the quantum anomaly but is related with the change of the Chern-Simons number of the $U(1)_{Y}$ field as $\Delta B=3\Delta N_{CS}[Y]$. Through this anomaly relation, the decay of each knot produces the baryon number $\Delta B\simeq\epsilon_{\mathrm{mix}}^{2}q_{1}q_{2}N_{\mathrm{link}}$. Therefore the linking number of the knot solitons can be a seed of the baryon number without leptogenesis.
In Refs. Kamada:2016eeb ; Kamada:2016cnb , baryogenesis is realized via decay of helical $U(1)_{Y}$ magnetic fields. While in their scenario the baryon number is produced during the EW phase transition (crossover), we now consider the case that the baryon number is produced after completion of the EW phase transition since $T_{rh}$ is lower than the EW scale, and hence it is not washed out by the sphaleron process.
In this scenario, we should consider carefully the creation of the knot solitons. Indeed, when the knot solitons are produced by the Kibble-Zurek mechanism, anti-knot solitons, in which the $\phi_{1}$ and $\phi_{2}$ strings are linked with the relatively opposite direction compared to knot solitons, are necessarily produced as well. This anti-knot has the Chern-Simons number with the opposite sign, $N_{CS}[A]\sim-\epsilon_{\mathrm{mix}}^{2}q_{1}q_{2}N_{\mathrm{link}}$. If there is no probability bias between the knots and anti-knots, their produced number should be equal, which results in that the net Chern-Simons number and hence the net baryon number are zero in total. To avoid this, one needs an effective chemical potential $\mu_{\mathrm{eff}}$ between them around the phase transition $T\sim T_{c}$. We assume that the asymmetry between them is given by the Boltzmann factor, leading to
$\left.n_{\mathrm{knot}}-\bar{n}_{\mathrm{knot}}\right|_{\text{production}}$ $\sim n_{\mathrm{knot}}\frac{\mu_{\mathrm{eff}}(T_{c})}{T_{c}}$ (24)
within the linear perturbation assuming $\mu_{\mathrm{eff}}(T_{c})/T_{c}\lesssim 1$.
Using this, the produced net baryon number from the decay of the (anti)knots is given as (assuming $N_{\mathrm{link}}\sim 1$)
$Y_{B}$ $\sim\left.\frac{n_{\mathrm{knot}}-\bar{n}_{\mathrm{knot}}}{s}\right|_{\text{decay}}\epsilon_{\mathrm{mix}}^{2}q_{1}q_{2}$
$\sim\left.n_{\mathrm{knot}}\right|_{\text{decay}}\,\frac{\mu_{\mathrm{eff}}(T_{c})}{T_{c}}\,\epsilon_{\mathrm{mix}}^{2}q_{1}q_{2}\frac{45}{2\pi}\frac{1}{g_{s}T_{rh}^{3}}$
$\sim 10^{-10}\frac{\mu_{\mathrm{eff}}(T_{c})}{T_{c}}\,\epsilon_{\mathrm{mix}}^{2}q_{1}q_{2}\left(\frac{T_{rh}}{100\,\mathrm{GeV}}\right)\left(\frac{10^{12}\,\mathrm{GeV}}{M}\right)\,.$ (25)
To get the observed baryon asymmetry, we need the mass of the knot soliton $M$ smaller than $10^{12}$ GeV. However, $M$ is typically given as $M\sim\mathcal{O}(10^{3})\,v/g$, and $v$ should satisfy $v\gtrsim 10^{8}$ GeV, in which the last inequality comes from astrophysical constraints on the axion decay constant, $f_{a}\gtrsim\mathcal{O}(10^{8})$ GeV Irastorza:2018dyq ; Chang:2018rso ; Hamaguchi:2018oqw ; Caputo:2024oqc . Thus successful baryogenesis works for $10^{8}\,\mathrm{GeV}\lesssim v\lesssim 10^{9}\,\mathrm{GeV}$.
We here comment on two possible origins of the chemical potential $\mu_{\mathrm{eff}}$. The first one is a new pseudo scalar $a^{\prime}$ that behaves as a time-dependent background. If $A_{\mu}$ couples to $a^{\prime}$ via $a^{\prime}F_{\mu\nu}\tilde{F}^{\mu\nu}$ in the Lagrangian, that term becomes $\partial_{t}a^{\prime}\epsilon^{ijk}A_{i}\partial_{j}A_{k}$, leading to $\mu_{\mathrm{eff}}\simeq\partial_{t}a^{\prime}N_{CS}[A]$. This resembles the Affleck-Dine mechanism Affleck:1984fy , Axiogenesis Co:2019wyp , and the production of helical magnetic fields from axion inflation Turner:1987bw ; Garretson:1992vt ; Anber:2006xt .
The second example of the chemical potential $\mu_{\mathrm{eff}}$ is the chiral charge asymmetry $n_{5}$, which can be produced from decays of heavy particles. For instance, the chiral charge of electrons can be produced Kamada:2018tcs ; Domcke:2020quw in a similar way to the GUT baryogenesis Yoshimura:1978ex ; Ignatiev:1978uf ; Weinberg:1979bt and is not washed out by the electron Yukawa coupling at high temperature. Through the chiral anomaly, this chiral asymmetry lowers (raises) energy of a helical (anti-helical) configuration of the $U(1)_{Y}$ gauge field, like chiral plasma instability Akamatsu:2013pjd . Thus, an imbalance between the knot and anti-knot might be induced as $\mu_{\mathrm{eff}}\simeq\mu_{5}N_{CS}[Y]$ with $\mu_{5}\sim n_{5}/T^{2}$.
In any case, we need some mechanism to give non-zero $\mu_{\mathrm{eff}}$ that leads to an asymmetry between knots and anti-knots. Obviously, it is required to have a closer look at the Kibble-Zurek mechanism in the presence of $\mu_{\mathrm{eff}}$ in order to obtain precise predictions of the asymmetry. After they are asymmetrically produced by the mechanism, the asymmetry is stored in the (anti-)knots as the net Chern-Simons number $N_{CS}[Y]$, and is converted into the baryon asymmetry by their decay. In this sense, the baryon asymmetry is “hidden” in the knot sector until the decay, and is eventually “released” into the Universe.
.4 Calculation of gravitational wave spectrum
We present formulae to calculate the GW spectrum from a cosmic string network. For more details, see Refs. Cui:2017ufi ; Cui:2018rwi ; Gouttenoire:2019kij . We assume that the network immediately approaches the scaling regime. The strings in the network move as Brownian motion and produce string loops by the reconnection. We model this in terms of velocity-dependent one-scale model Martins:1995tg ; Martins:1996jp ; Martins:2000cs with a loop chopping efficiency $\bar{c}=0.23$. Then, the produced number density of the loops per time at cosmic time $t_{i}$ is given by
$\frac{dn_{\mathrm{loop}}}{dt_{i}}=C_{\mathrm{eff}}(t_{i})t_{i}^{-4},$ (26)
with an initial length $l=\alpha t_{i}$. We take $\alpha\simeq 0.1$ and $C_{\mathrm{eff}}(t_{i})=0.41$ ($5.5$) for the matter dominated (radiation dominated) Universe.
After the production, the loops oscillate and shrink by emitting GW. The length at time $t$ is given as
$l=\alpha t_{i}-\Gamma G\mu(t-t_{i})$ (27)
with the dimensionless constant $\Gamma\simeq 50$ Vilenkin:1981bx ; Turok:1984cn ; Quashnock:1990wv ; Blanco-Pillado:2013qja ; Blanco-Pillado:2017oxo . Each oscillation mode $k$ radiates GW with the frequency $f_{\mathrm{emit}}=2k/l$ ($k=1,2,\cdots$). After emission at time $\tilde{t}$, the frequency redshifts due to the cosmic expansion until today as
$f=\frac{a(\tilde{t})}{a(t_{0})}\frac{2k}{l(\tilde{t})}=\frac{a(\tilde{t})}{a(t_{0})}\frac{2k}{\alpha t_{i}-\Gamma G\mu(\tilde{t}-t_{i})}\,,$ (28)
where $t_{0}$ is the current time.
Summing over all oscillation modes, the energy density of GW per unit frequency observed today is given as
$\Omega_{\mathrm{GW}}(f)=\frac{f}{\rho_{c}}\frac{d\rho_{\mathrm{GW}}}{df}=\sum_{k}\Omega_{\mathrm{GW}}^{(k)}(f)$ (29)
with
$\Omega_{\mathrm{GW}}^{(k)}(f)$ $=\frac{1}{\rho_{c}}\frac{2k}{f}\frac{\Gamma_{k}G\mu^{2}}{\alpha+\Gamma G\mu}$
×$\times\int_{t_{F}}^{t_{0}}d\tilde{t}\,\frac{C_{\mathrm{eff}}(t_{i})}{t_{i}^{4}}\left(\frac{a(\tilde{t})}{a(t_{0})}\right)^{5}\left(\frac{a(t_{i})}{a(\tilde{t})}\right)^{3}\Theta(t_{i}-t_{F})\,,$ (30)
where $\rho_{c}=3H_{0}^{2}/8\pi G$ is the critical density and $\Gamma_{k}=\Gamma k^{-4/3}/3.6$ Blanco-Pillado:2013qja ; Blanco-Pillado:2017oxo . $t_{F}$ is the formation time of the string network, which can be taken to be 00 essentially. We assume that the GW emission is dominated by cusps. The relation between $\tilde{t}$ and $t_{i}$ is given by Eq. (28). $\Theta$ is the Heaviside step function.
We assume sudden transitions between the radiation dominated and knot dominated eras. The summation over the modes $k$ is performed by the method presented in Ref. Blasi:2020mfx , to be taken up to $k_{\mathrm{max}}=10^{12}$, which gives the slopes $f^{-1/3}$ Blasi:2020wpy ; Blasi:2020mfx at higher frequency regime of the spectra with the knot dominated era. Although one may truncate the summation within smaller values, it affects only an asymptotic behavior at high frequency. Fig. 4 is obtained by plotting Eq. (29) utilizing these techniques.